OpenStudy (anonymous):

Solving systems of equations algebraically and verifying the solution using matrices or graphically. Please help :)

OpenStudy (anonymous):

8x-2y=5 -12x+3y=7

OpenStudy (anonymous):

could you possibly help?

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

how would that look sir?

OpenStudy (midhun.madhu1987):

This equations have No solutions

OpenStudy (anonymous):

really? this is one on a practice paper, usually it would say that it doesn't have a solution

OpenStudy (anonymous):

my brain is fried ive been doing this paper for the past 2 hours

OpenStudy (midhun.madhu1987):

 See, both their slopes are equal and their y-intercepts are different. so they dont have a solution in common...